# Security and Quality in LLM-Generated Code: A Multi-Language, Multi-Model Analysis

### v1

@@ Abstract
Artificial Intelligence (AI)-driven code generation tools are increasingly used throughout the software development lifecycle to accelerate coding tasks. However, the security of AI-generated code using Large Language Models (LLMs) remains underexplored, with studies revealing various risks and weaknesses. This paper analyzes the security of code generated by LLMs across different programming languages. We introduce a dataset of 200 tasks grouped into six categories to evaluate the performance of LLMs in generating secure and maintainable code. Our research shows that while LLMs can automate code creation, their security effectiveness varies by language. Many models fail to utilize modern security features in recent compiler and toolkit updates, such as Java 17. Moreover, outdated methods are still commonly used, particularly in C++. This highlights the need for advancing LLMs to enhance security and quality while incorporating emerging best practices in programming languages.
@@ 1 Introduction
The rapid advancement of AI technologies has led to the widespread use of LLMs in generating source code for various programming tasks. LLMs have performed remarkably in various natural language processing and generation tasks since a significant turning point with transformer models [1]. Different LLMs have emerged as powerful resources for developers since these tools can generate functional code across multiple programming languages.
@@
In 2018, Microsoft Visual Studio released the IntelliCode extension [2], which offers AI-powered development features that provide limited insights by analyzing code context using machine learning. In 2021, GitHub introduced Copilot, an AI-driven code assistant designed to improve coding quality by training on extensive real-world code repositories [3], enabling it to provide coding recommendations across various programming languages and frameworks. Since GitHub introduced Copilot, AI-driven code generation adoption has grown in the software development lifecycle [4] where complex models are used to perform specific tasks like writing code for software engineers [5].
@@
One major challenge for AI-driven code generation is ensuring code quality, where key metrics include validity, correctness, security, reliability, and maintainability. Current research stresses the importance of ensuring security in LLM-generated code [6, 7]. Several other studies have shown that LLMs are capable of producing code that is both secure and vulnerable [8, 6, 5]. The factors affecting the security properties of the code generated by these models are not identified. Multiple studies highlight the need for better user guidelines and awareness when interacting with AI tools [6, 7, 9] to enhance the quality properties of generated code. The security quality of the code generated for identical scenarios may differ depending on the chosen programming language, which is considered an explainable issue in artificial intelligence [10, 7].
@@
This study addresses the gap in the literature by investigating the factors that impact the security of code produced by LLMs, with a focus on the influence of programming language selection. Our contributions include the development of a comprehensive dataset designed to evaluate AI-generated code across multiple domains, such as problem-solving, algorithms, and other key areas. Additionally, we evaluate the ability of various LLMs (claude-3.5, gemini-1.5, codestral, GPT-4o, llama-3) to generate secure and functional code in different languages (Python, Java, C++, and C). Moreover, we conduct a security analysis of AI-generated code using static security analysis tools (SAST) to identify common vulnerabilities and weaknesses. Lastly, this research conducts a comparative study of the security properties of AI-generated code across different programming languages, highlighting the main strengths and limitations of each AI tool in contexts such as semantic correctness and security.
Contributions. We make the following contributions. (1) New dataset for LLM-based coding. We introduce a new manually vetted dataset of 200 programming tasks classified into six categories that can be used by the research community for evaluating the performance of LLMs. (2) Comprehensive analysis LLM-generated codes. We comprehensively explore the quality attributes and security of the code generated by LLM under the same evaluation condition and using the proposed dataset. (3) In-depth security analysis. We conduct a comprehensive study in different languages that highlight characteristics, issues, and language-specific differences in security vulnerabilities with LLMs used to generate code.
@@ 2 Related Work
Several works explored the use of LLMs in a variety of software development activities [23, 6, 13, 4, 16, 14, 7, 9, 5, 8, 22, 19, 24, 25, 26]. This section examines key studies, emphasizing their methods, applications, and the analyzed features. The comparative overview helps place our work in the broader research context.
Huang et al. [23] reviewed pre-trained models and LLMs for generative tasks in software engineering, highlighting models like BERT, general transformers, and ChatGPT. They categorized tasks into requirements generation, code generation, test case generation, patch generation, optimization, summarization, and code translation. Our research expands on this by examining the effectiveness of LLMs, specifically in code generation, enhancing the understanding of their strengths and limitations in this area.
Perry et al. [6] investigated the security of LLM-based codes, finding that such codes often had more security flaws than human-written codes, with LLMs displaying overconfidence in code security. Sandoval et al. [13] observed a 10% increase in vulnerabilities in LLM-based C programming. Asare et al. [4] found that GitHub Copilot could enhance security for complex problems but had minimal impact on simpler tasks. These studies highlight the need for caution when using AI tools. Our research expands by assessing four LLMs across additional programming languages, C++ and Java, for further evaluation.
Yetiştiren et al. [16] assessed the quality of code produced by three AI code assistants using the HumanEval benchmark dataset [17], finding correctness rates of 65.2%, 46.3%, and 31.1% for ChatGPT, GitHub Copilot, and Amazon CodeWhisperer, respectively. Factors like function names, input, and descriptions were evaluated, and SonarQube [27] was used to assess security, maintainability, and reliability. All three tools were found capable of generating secure code. In contrast, our study expands to four languages and incorporates a broader range of dataset scenarios.
@@
Nair et al. [9] explored ChatGPT’s effectiveness in generating hardware code using Common Vulnerability Enumerations (CWE-1194), demonstrating the possibility of guiding AI to avoid common security flaws. Elgedawy et al. [5] analyzed GPT-3.5, GPT-4, Bard, and Gemini, showing that using security personas reduced vulnerabilities, especially in GPT-3.5, GPT-4, and Bard. Siddiq et al. [8] introduced the SALLMS framework to evaluate LLMs systematically for security. Schuster et al. [22] and Wu et al. [19] highlighted challenges, such as data poisoning and LLMs’ reduced effectiveness in handling complex security issues, emphasizing the need for stronger defenses and better vulnerability detection.
@@
Table I summarizes related work based on the programming languages used in LLM-generated code, the LLM(s) employed, and a summary of the prompt descriptions. Table II outlines related work by the quality characteristics addressed and the security analysis methods used to evaluate the generated code. Additionally, the table provides the overall impact of AI code generation on code quality based on the respective study’s experiments and results.
@@
Our Work. Most existing research focuses on exploring LLM code generation behavior but lacks in-depth analysis of factors affecting the quality and security of the generated code, beyond user demographics and seniority level. In contrast, our work advances the literature by examining the relationship between programming language features and the quality of LLM-generated code. We evaluate the validity, correctness, security, maintainability, consistency, intentionality, adaptability, and responsibility of the code across four programming languages and five LLMs.
@@
Table II: A summary of the related work, highlighting the features and methods used to evaluate the quality of generated code. Overall Impact (OI): Negative (N), Positive (P), and Negative impact due to the use of an inappropriate dataset for testing code security (N*). Attributes: Validity, Correctness, Security, Reliability, and Maintainability. Methods: Manual, Static Scan, and Runtime Scan.
@@ 3 Methodology
3.1 LLMs Selection
The choice of LLMs is based on several metrics, such as popularity, user base, reputation, support of diverse programming languages, and performance (accuracy, efficiency) in the code generation process. These metrics are derived from academic literature and industry benchmarks, ensuring that our study is representative. The LLMs selected in this study, Table III, cover a wide range, each with distinct strengths in terms of efficiency in the code generation process, language support, and overall accuracy. A significant factor in our selection is the diversity of their underlying LLM architecture, whereas they vary in the context of parameter size, training datasets, and decoding strategies. These differences might directly influence the security and quality of the generated code, making a comparative analysis of its output important to understand the strengths and weaknesses of each of these models.
@@ 3.1 LLMs Selection
  Provider Ref Model Short
  OpenAI [28] GPT-4o GPT-4o
@@
CLAUDE [30] claude-3-5-sonnet-20240620 claude-3.5
Mistral [31] codestral-2405 codestral
@@ 3.2 Programming Languages Selection
Each programming language possesses distinct attributes that influence security outcomes in code generation (i.e., Python’s dynamic typing versus C++’s static typing can introduce different vulnerabilities and bugs). Java and Python benefit from automatic memory management, reducing memory-related errors compared to the manual memory management required in C and C++. Thus, the selected languages provide a comprehensive view of how LLMs address these differences. Additionally, this selection ensures that the study’s findings are applicable to a wider range of real-world software development scenarios, as these languages rank among the top five most used in 2024, making the research highly relevant to many users.
@@ 3.3 Dataset
To thoroughly evaluate the performance of LLMs, we curated a set of 200 prompt descriptions aimed at testing multiple code generation facets. The selection of tasks was undertaken to ensure comprehensive coverage of key programming paradigms and secure coding practices. The dataset is detailed in section 4.
@@ 3.4 Environment Setup
To ensure consistency and reproducibility, we created a standardized environment for generating, compiling, executing, and validating code. Designed for diverse programming languages and tasks, it supports multiple languages and unit test execution. All experiments were conducted on a Lenovo ThinkPad E570 with a 7th-gen Intel® Core™ i7, 16 GB DDR4 RAM, and a 256 GB SSD. This hardware was chosen for its availability, efficiency, and portability, making it ideal for language model integration, code generation, and multi-language compilation. Debian 12 was selected for its stability, efficient package management, and minimal resource usage, providing a reliable platform for cross-language development and testing. The following software packages and versions were used:
  1. Java. To handle Java code compilation and execution, we used the long-term support (LTS) Java version (OpenJDK version 17.0.8). This version supports the latest features of the Java language, ensuring compatibility with modern programming constructs and practices generated by the LLMs.
  2. Python. For Python code, we utilized Python version 3.11.9. This version was chosen for its compatibility with the latest Python libraries and features, ensuring that the generated Python code was evaluated in an up-to-date runtime environment.
  3. C and C++. The compilation of the C and C++ codes was carried out using CMake version 3.28.6. For C++ specifically, we configured the CMAKE_CXX_STANDARD to version 17, which gained popularity after its release in 2017 [34], ensuring that all C++ code generated by the models adhered to the C++17 standard. This choice was made to support modern C++ features, such as structured bindings and inline variables, which are common in LLM-generated code.
@@ 3.5 LLM Integration and Code Generation
Based on the LLMs selected in Table III, each model was used to generate responses in the four selected programming languages. To streamline the code generation process, the LLMs were interfaced with a custom written Python program that methodically dispatched task prompts in sequential order to each model and collected the resultant code output. Upon code generation, another custom-written Python program was used to systematically parse and arrange the results within a structured file system, assigning appropriate extensions pertinent to the programming language (i.e., .py for Python, .java for Java). Each file was named according to the task prompt identifier, the LLM model used, and the language, facilitating traceability and ease of comparison.
4,000 code files were generated, comprising 200 files per language for each LLM. This dataset formed the foundation for the subsequent evaluation procedures, ensuring that each LLM was evaluated using an identical set of tasks and programming languages.
@@ 3.6 Quality Evaluation
To evaluate the quality of the generated code, analysis is performed based on several quality metrics, including syntax validity, functional correctness, code lines, reliability, maintainability, and security [16, 8]. We use two types of evaluation methods: ① Manual evaluation by a human expert, where two developers participated in the semantic evaluation and wrote the unit testing files; ② Automatic static secure code scanning using the SonarQube [27], a static code scanning tool. Although human evaluation metrics may not scale, they are still considered the golden standard for evaluating the output of LLMs and NLP tasks. Using automatic analysis tools, we will use human metrics to evaluate a small-scale set of LLM examples and generalize beyond the small sample of human evaluators. A review of recent literature that identifies these evaluation methods as an important approach in code quality evaluation supports the selection of the metrics and tools mentioned above.
3.6.1 Manual Method
Two developers, with two and ten years of software development experience, are hired to participate in subsequent evaluations. In addition, the first and second authors participated in the process of progress coordination and reviews.

### v2

@@ Abstract
Artificial Intelligence (AI) driven code generation tools are increasingly used throughout the software development lifecycle to accelerate coding tasks. However, the security of AI-generated code using large language models (LLMs) remains underexplored, and recent studies have revealed various risks and weaknesses. This paper presents a measurement study of LLM-generated code across four programming languages (Python, Java, C++, and C) and five widely used LLM families. We construct a manually curated dataset of 200 programming tasks, grouped into seven functional and security-relevant categories, each with language-neutral specifications. For every combination of task, language, and model, we generate code and evaluate it along three axes: syntactic validity and compilation success, semantic correctness using 4,000 per program unit test files, and software quality and security using SonarQube and CodeQL, complemented by manual review of key static analysis findings. Our results show clear language effects: Python and Java achieve higher compilation and semantic correctness rates and produce fewer security findings than C and C++, where we observe more memory safety issues, hard-coded secrets, and cryptographic misuses. We also find that many models fail to make use of modern security features available in recent compiler and toolkit updates (i.e., in Java 17), and that outdated methods remain common, particularly in C++. These findings highlight the need to advance LLMs so that they better align with emerging secure coding practices and language-specific best practices. All code and data are available at GitHub.
@@ 1 Introduction
The rapid advancement of artificial intelligence (AI) technologies has driven the widespread adoption of large language models (LLMs) for generating source code across a broad range of programming tasks. In particular, LLMs have demonstrated remarkable performance in natural language understanding and generation. Building on these capabilities, recent LLMs have emerged as powerful tools for software developers, enabling the automated generation of functional code in multiple programming languages.
@@
In 2018, Microsoft Visual Studio released the IntelliCode extension [1], which offers AI-powered development features that provide limited insights by analyzing code context using machine learning. In 2021, GitHub introduced Copilot, an AI-driven code assistant designed to improve coding quality by training on extensive real-world code repositories [2], enabling it to provide coding recommendations across various programming languages and frameworks. Since GitHub introduced Copilot, AI-driven code generation adoption has grown in the software development lifecycle [3], where complex models are used to perform specific tasks like writing code for software engineers [4].
@@
One major challenge for AI-driven code generation is ensuring code quality, where key metrics include validity, correctness, security, reliability, and maintainability. Current research stresses the importance of ensuring security in LLM-generated code [5, 6]. Several other studies have shown that LLMs are capable of producing code that is both secure and vulnerable [7, 5, 4]. The factors affecting the security of the code generated by these models are not identified. Multiple studies highlight the need for better user guidelines and awareness when interacting with AI tools [5, 6, 8] to enhance the quality properties of generated code. The security quality of the code generated for identical scenarios may differ depending on the chosen programming language, which is considered an explainable issue [9, 6].
@@
This work addresses this gap by investigating the factors that impact the security of LLMs-generated code, with a focus on the influence of language selection. Our contributions include the development of a dataset designed to evaluate AI-generated code across multiple domains, such as problem-solving, algorithms, and other key areas. Additionally, we evaluate the ability of various LLMs (claude-3.5, gemini-1.5, codestral, GPT-4o, llama-3) to generate secure and functional code in different languages (Python, Java, C++, and C). Moreover, We conduct a quality and security analysis of LLM-generated code using static security analysis tools (SAST) to identify common vulnerabilities and weaknesses, and quantify their distribution across models and languages, focusing on how the known CWE categories manifest in a unified evaluation setting. Lastly, this research conducts a comparative study of the security properties of AI-generated code across different programming languages, highlighting the main strengths and limitations of each AI tool in contexts such as semantic correctness and security.
Research Questions. We address this problem through the following research questions. RQ1. How does the choice of programming language affect the security of code generated by large language models (LLMs)? RQ2. How do different LLMs compare in generating syntactically valid and semantically correct code across multiple programming languages? RQ3. Which security vulnerabilities are most prevalent in LLM-generated code, and how do they vary across models and programming languages?
Contributions. We make the following contributions. (1) New dataset for LLM-based coding. We introduce a new manually vetted dataset of 200 programming tasks from seven categories that can be used by the research community for evaluating the performance of LLMs. (2) Comparative evaluation of LLM-generated code. We evaluate the quality and security of LLM-generated code across multiple programming languages and models using a unified dataset and a consistent evaluation setup. This enables direct cross-language and cross-model comparisons and exposes differences that are missed in single-language or single-model studies. We treat code quality and security as complementary dimensions, as low-quality code often correlates with security vulnerabilities and secure coding practices align with readability and maintainability. Joint evaluation establishes a baseline for analyzing their interactions and trade-offs in LLM-generated code. (3) In-depth security analysis. We analyze LLM-generated code across multiple programming languages to identify language-specific characteristics, recurring issues, and distinct security vulnerability patterns.
@@
Ullah et al. [21] 2024 200 tasks
Siddiq et al. [22] 2022 SecurityEval dataset [19]
Khare et al. [23] 2024 25 CWEs, dataflow based prompt
Asare et al. [asare2024llms] 2024 200 Tasks
Tony et al. [24] 2023 LLMSecEval dataset [19]
Lenarduzzi et al. [25] 2021 LLMSecEval dataset [19]
@@ 2 Related Work
Several works explored LLMs’s use in a variety of software development activities [26, 5, 11, 3, 14, 12, 6, 8, 4, 7, 20, 17, 21, 27, 22, 24, 25]. This section examines key studies, emphasizing their methods, applications, and the analyzed features. The comparative overview helps place our work in the broader research context.
Huang et al. [26] surveyed pre-trained models and LLMs for software engineering tasks, including requirements, code, test case, patch generation, optimization, summarization, and code translation, covering models such as BERT, general transformers, and ChatGPT. Our work builds on this line of research by focusing specifically on code generation and evaluating the effectiveness of LLMs in this setting. Perry et al. [5] analyzed the security of LLM-generated code and showed that it often contains more security flaws than human-written code, with LLMs exhibiting overconfidence in code safety. Sandoval et al. [11] reported a 10% increase in vulnerabilities in LLM-generated C code. Asare et al. [3] found that GitHub Copilot can improve security for complex tasks but provides limited benefit for simpler ones. In contrast to prior work, we evaluate four LLMs across multiple programming languages, including C++ and Java, to enable broader cross-language and cross-model analysis.
Yetiştiren et al. [14] evaluated the quality of code generated by three code assistants using the HumanEval benchmark [15]. They reported correctness rates of 65.2%, 46.3%, and 31.1% for ChatGPT, Copilot, and CodeWhisperer, respectively, and analyzed the impact of function names, inputs, and problem descriptions. Security, maintainability, and reliability were assessed using SonarQube [28], with all tools producing code deemed secure. Our work differs by evaluating four programming languages and using a broader and more diverse set of dataset scenarios.
@@
Nair et al. [8] explored ChatGPT’s effectiveness in generating hardware code using Common Vulnerability Enumerations (CWE-1194), showing the possibility of guiding AI to avoid common flaws. Elgedawy et al. [4] analyzed GPT-3.5, GPT-4, Bard, and Gemini, showing that using security personas reduced vulnerabilities, especially in GPT-3.5, GPT-4, and Bard. Siddiq et al. [7] introduced the SALLMS framework to evaluate LLMs systematically for security. Schuster et al. [20] and Wu et al. [17] highlighted challenges, such as data poisoning and the reduced effectiveness of LLMs in handling complex security issues, highlighting the need for stronger defenses and better vulnerability detection.
Ullah et al. [21] evaluated modern LLMs for vulnerability detection and identified limitations in consistency, reasoning accuracy, and robustness, particularly under minor code changes and real-world conditions. Pearce et al. [27] analyzed the security implications of code generated by GitHub Copilot and showed that the tool frequently produces vulnerable patterns, especially when prompts are underspecified or ambiguous. Siddiq and Santos [22] introduced SecurityEval, a dataset of 130 Python programs mapped to 75 common weakness enumerations (CWEs), to evaluate the security of machine-generated code. Using SecurityEval, they assessed tools such as GitHub Copilot and InCoder and found that these models often generate code with security vulnerabilities, emphasizing the need to evaluate generated code beyond functional correctness to include security.
Asare et al. [12] evaluated the ability of LLMs to detect and reason about security vulnerabilities in source code across multiple languages and tasks. The study showed that most LLMs fail to identify deeper security flaws, even when the code is syntactically correct, highlighting the need for more reliable, security-focused benchmarks and evaluation methods. Tony et al. [24] introduced LLMSecEval, a dataset of natural language prompts designed to assess LLM performance on security-related questions across common software security topics, enabling consistent evaluation of security understanding. Together, these studies stress the importance of evaluating LLMs beyond code generation, with explicit focus on security reasoning.
Lenarduzzi et al. [25] compared six code analysis tools and evaluated their issue detection effectiveness, result consistency, and precision and reported substantial differences in detection coverage and overlap among tools, showing that reliance on a single tool can miss vulnerabilities or introduce false positives. Khare et al. [23] performed a large-scale evaluation of LLM-based vulnerability detection, testing 16 models on 5,000 code samples from five datasets of C, C++, and Java. The authors proposed dataflow-inspired prompting strategies and showed that LLMs can outperform static and deep learning-based tools for specific vulnerability classes while providing interpretable explanations.
@@
Table I summarizes prior work by language, LLM, and prompt for LLM-generated code. Table II categorizes related work by the quality attributes examined and the security analysis methods applied to evaluate generated code, and reports the overall impact of AI code generation on code quality based on each study’s experimental results.
@@
Our Work. Most prior studies analyze LLM code generation behavior but provide limited insight into the factors affecting code quality and security beyond user demographics and experience levels. This work examines how programming language characteristics influence the quality of LLM-generated code. We evaluate validity, correctness, security, maintainability, consistency, intentionality, adaptability, and responsibility across four languages and five LLMs. This study is a quantitative measurement study: it introduces no new vulnerability classes or analysis techniques, but systematically compares the behavior of existing LLMs across languages and tasks under a unified setup.
@@
Table II: Summary of related work. Overall Impact (OI): Negative (N), Positive (P), and Negative due to the use of an inappropriate dataset for security (N*). Attributes: Validity, Correctness, Security, Reliability, and Maintainability. Methods: Manual, Static Scan, and Runtime Scan.
@@ V C Se Re Ma Pe M S R
Khare et al. [23] P
Asare et al. [3] N
Tony et al. [24] N
Lenarduzzi et al. [25] N
Ullah et al. [21] N
Siddiq et al. [7] N
@@ 3 Methodology
We selected LLMs using a two-step process. First, we defined eligibility criteria based on prior literature and industry benchmarks: ① widespread adoption in software engineering practice, ② demonstrated performance in program synthesis, reasoning, and security-related tasks, ③ official support for code generation in Python, Java, C++, and C, ④ stable and accessible APIs during data collection, and ⑤ architectural diversity to avoid evaluating closely related model variants. Second, we applied these criteria to available models and list the selected LLMs in Table III. The selected models are widely used, frequently evaluated in prior studies, support all target languages through stable APIs, and represent distinct training and design choices, enabling systematic comparative analysis.
@@
  Provider Ref Model Version Short
  OpenAI [29] GPT-4o-202407 GPT-4o
@@
Anthropic [31] claude-3-5-sonnet-20240620 claude-3.5
Mistral AI [32] codestral-2405 codestral
@@ 3.1 Programming Languages Selection
Each programming language exhibits attributes that affect security outcomes in code generation (e.g., Python’s dynamic typing versus C++’s static typing leads to different classes of bugs and vulnerabilities). Java and Python provide automatic memory management, reducing memory-related errors relative to C and C++, which rely on manual memory management. These languages therefore enable analysis of how LLMs handle language-specific security properties. The selection also improves external validity, as these languages rank among the top five most used in 2024 and reflect common real-world development settings.
@@ 3.2 Dataset
To comprehensively evaluate LLM performance, we curated 200 prompt descriptions targeting diverse aspects of code generation. Task selection ensured broad coverage of core programming paradigms and secure coding practices. Details are provided in section 4.
@@ 3.3 Environment Setup
To ensure consistency and reproducibility, we used a standardized environment for code generation, compilation, execution, and validation across tasks and programming languages, including unit test support. All experiments ran on a Lenovo ThinkPad E570 with a 7th-generation Intel® Core i7 processor, 16 GB DDR4 RAM, and a 256 GB SSD. This configuration provides sufficient performance and portability for LLM integration, code generation, and multi-language compilation. We used Debian 12 for its stability, package management, and low resource overhead, ensuring a consistent platform for cross-language testing. The following software packages and versions were used. ❶ Java. We compiled and executed Java code using OpenJDK 17.0.8 (LTS), which supports modern Java language features and ensures compatibility with constructs generated by LLMs. ❷ Python. We evaluated Python code using Python 3.11.9 to ensure compatibility with current features and libraries. ❸ C and C++. We compiled C and C++ code using CMake 3.28.6. For C++, we set CMAKE_CXX_STANDARD to 17, a widely adopted standard since its 2017 release [35], ensuring support for modern C++ features produced by LLMs.
@@ 3.4 LLM Integration and Code Generation
Based on the LLMs listed in Table III, we generated code for each task in four programming languages. A custom Python script sequentially dispatched task prompts to each model and collected the generated code. All inputs used the original task descriptions without prompt engineering. After generation, a second Python script parsed and organized the outputs into a structured file system, assigning language-specific file extensions (e.g., .py for Python and .java for Java). Each file name encoded the task identifier, LLM, and programming language to support traceability and comparison. The first author verified the total number of generated programs to ensure complete collection.
As a result, 4,000 code files were generated, comprising 200 files per language for each LLM. This dataset formed the foundation for the subsequent evaluation procedures, ensuring that each LLM was evaluated using an identical set of tasks and programming languages.
@@ 3.5 Quality Evaluation
To evaluate code quality, we analyze syntax validity, functional correctness, code length, reliability, maintainability, and security [14, 7] using two evaluation approaches. ① Manual evaluation by human experts where two software engineers with four and ten years of professional experience participated in the evaluation, with the first and second authors coordinating and reviewing the process. Drafting, validating, and executing test cases to assess semantic correctness required over 800 person-hours across five months. ② Automated static analysis using SonarQube [28] and CodeQL [36], which support all target languages. Automated tools provide scalable coverage of quality and security metrics, but have known limitations when used in isolation. Human evaluation remains the gold standard for assessing LLM outputs. We therefore complement automated analysis with human evaluation on a representative subset of generated programs to support generalization beyond a limited number of evaluators. Prior work identifies these methods as effective for code quality and security assessment, supporting our choice of metrics and tools.
3.5.1 Semantic Soundness